OpenStudy (anonymous):

y varies inversely as three times x. When x = 6, y = 2. Find x when y = 12.

OpenStudy (saifoo.khan):

i did that already.

OpenStudy (anonymous):

y = (1/3) * x. that means that x = 3y. for x=6, y=2. ( 2*3 = 6 ) if y = 12, then x = 3 * 12 = 36.